OpenStudy (anonymous):

Please help solve this equation. sin x cos x = √3 / 4

OpenStudy (jdoe0001):

$$ sin(x)cos(x)=\cfrac{\sqrt{3}}{4} \implies sin(x)cos(x)=\cfrac{\sqrt{3}}{2\times2}\\ 2sin(x)cos(x)=\cfrac{\sqrt{3}}{2} \&\& \ \color{blue}{ 2sin(x)cos(x)=sin(2x)}\\ sin(2x)=\cfrac{\sqrt{3}}{2} \implies sin^{-1}(sin(2x))=sin^{-1}\pmatrix{\cfrac{\sqrt{3}}{2}} $$

OpenStudy (jdoe0001):

can you take it from there?

OpenStudy (jdoe0001):

@jigglygurl2 still need help with that?